OpenStudy (keana):

find the value of k that would make the left side of the equation a perfect square trinomial: x^2+kx+25=0

OpenStudy (anonymous):

Ok let me see

OpenStudy (anonymous):

Do you have choices?

OpenStudy (keana):

no

OpenStudy (anonymous):

I would say K will equal 30

OpenStudy (keana):

can you explain it? :)

OpenStudy (anonymous):

Ok wait up

OpenStudy (anonymous):

Take the square root of the first and last terms..... 3x and 5 Multiply those together..... 15x Double that answer and that is what your middle term needs to be if you are going after a perfect square trinomial. (Your given polynomial already has the "x" part so you are only looking for the 30.) I had the same question so this is what i put

OpenStudy (raden):

x^2+kx+25=0 see the expanding of (x + 5)^2 = x^2 + 10x + 25 now just identicaly with the equation above, you will see the value of k there